# A Decade of Privacy-Relevant Android App Reviews: Large Scale Trends

### 04 Mar 2024 v1

@@ Abstract
We present an analysis of 12 million instances of privacy-relevant reviews publicly visible on the Google Play Store that span a 10 year period. By leveraging state of the art NLP techniques, we can examine what users have been writing about privacy along multiple dimensions: time, countries, app types, diverse privacy topics, and even across a spectrum of emotions. We find consistent growth of privacy-relevant reviews, and explore topics that are trending (such as Data Deletion and Data Theft), as well as those on the decline (such as privacy-relevant reviews on sensitive permissions). We find that although privacy reviews come from more than 200 countries, 33 countries provide 90% of privacy reviews. We conduct a comparison across countries by examining the distribution of privacy topics a country’s users write about, and find that geographic proximity is not a reliable indicator that nearby countries have similar privacy perspectives. We uncover some countries with unique patterns and explore those herein. Surprisingly, we uncover that it is not uncommon for reviews that discuss privacy to be positive (32%); many users express pleasure about privacy features within apps or privacy-focused apps. We also uncover some unexpected behaviors, such as the use of reviews to deliver privacy disclaimers to developers. Finally, we demonstrate the value of analyzing app reviews with our approach as a complement to existing methods for understanding users’ perspectives about privacy.
@@ 1 Introduction
Getting input about users’ perspectives on privacy issues from millions of user reviews from hundreds of countries has not been feasible until recently. Advances in natural language processing (NLP) and large language models (LLMs) permit complex analysis of enormous text corpora. Here, we leverage these advances to investigate users’ privacy opinions and concerns from a novel perspective: we study 12.3M (million) privacy-related reviews, extracted from $\sim$2B (billion) public reviews on Google Play spanning 10 years (Jan. 2013–Feb. 2023). These privacy reviews come from more than 200 countries or regions and 160K (thousand) apps that span every Play app category.
@@ 3.1 Text Analysis Pipeline
We build our analysis pipeline to reproduce the components in Hark, an end-to-end system for retrieval and analysis of privacy-related feedback leveraging state-of-the-art techniques in NLP [30]. An overview of our data-analysis pipeline is in Figure 1. First, the privacy classifier identifies the privacy-related feedback from unstructured text. The issue generation model takes in this privacy feedback and dynamically generates meaningful, fine-grained issues (covering both known and newly emerging issues) describing the privacy aspects discussed within each review text. The theme creation component groups these issues into thematic clusters and assigns a succinct title to each. Our data-analysis pipeline also includes an emotion classifier that dissects each review’s text across 28 emotions (e.g., anger, fear, joy and confusion).
@@
We obtained models and training data from Hark [30] and reproduced most of the pipeline, with modifications to the privacy classifier to improve the breadth of topics identified. First, we expand the privacy taxonomy from the original paper (25 concepts) to include 89 privacy-relevant concepts(see Appendix D). The prior taxonomy already included concepts from multiple known taxonomies. However, when we asked five privacy experts (with at least 5 years of academic/industry experience in privacy research and development) to manually examine the original taxonomy [30], they observed many missing topics. For example, we added Data Hiding, Opt-out and Location concepts, and added specificity to the Anonymity and Personal Data types.
@@
Our classifier has a 0.95 ROC AUC, 87% precision, and 86% recall on our new diversified test set [^1] We set a 0.8 prediction score threshold for best accuracy and verified results qualitatively., and a similar ROC AUC of 0.88 (vs 0.92) on the original Hark test set [30]. In comparison, the privacy classifier from [30] has 0.87 ROC AUC, 89% precision, and 51% recall on our new test set, demonstrating that our classifier performs better at capturing the diversity of our taxonomy. We run additional experiments with a range of model architectures. See Appendix B for details on other baselines evaluated. We also performed qualitative assessment of all 50 false positives/negatives of our classifier, noticing primarily issues with reviews that were short ($\leq$ 10 words) and ambiguous (missing context, causing multiple possible interpretations). For example, a short review “Taking too much data” was labeled as ‘privacy,’ when the privacy expert (based on other reviews seen) interpreted it to be a complaint about the app consuming the limited mobile data bandwidth and annotated it as ‘not-privacy’; whereas an ambiguous review “msgs have been deleted, but for some reason they remain in place” was labeled as ‘not-privacy’ (could be seen as a bug in app functionality), when the privacy expert interpreted it as data deletion control not working as expected (a privacy concern). 13 of the 50 false positives/negatives were short; and the rest (though longer) were still ambiguous. Despite missing these short and ambiguous cases, our classifier had acceptable performance.
@@ 3.2 Data Description
Our initial dataset consists of all $\sim$2B publicly visible Google Play app reviews for apps with $>$10K installs, spanning 10 years from Jan 2013–Feb 2023. The dataset was anonymized – without any user identifying information, and sanitized – any detected fake/spam reviews were removed. Each review is associated with the review text, its language, country, submission time, star rating, the corresponding app’s package name, the app’s developer-specified category information and (if available) finer-grained app-type information (such as ‘Rideshare & Taxis’ for a rideshare app, which belongs to the ‘Maps & Navigation’ developer-specified category). Note that we do not generally use developer-specified categories as they are too broad [66]. Instead we use the more specific app types that are displayed on Play when viewing apps.
In our dataset, $\sim$65of reviews are in non-English languages. Since our classifiers work on English texts, we leverage Google’s Translation API to translate reviews in 24 non-English languages to English (Appendix A[4]. The final set of English reviews (including translations) constitutes 98% of the initial dataset and contains 1.9B reviews for 160K apps. These reviews come from all developer-specified categories and 445 app types, with representation from more than 200 countries and territories in the world.
@@ 3.5 Ethical considerations
Our work uncovered reviews that contained ads for spying services. We disclosed our findings to Google Play and ensured these harmful reviews were removed before publication.
@@ 3.6 Limitations
Our work focuses on Android users; however, other sizeable platforms exist, raising generalizability concerns depending on the ratio of android users in a country. We note that Android is the most popular mobile operating system in the world and tends to be even more popular in non-WEIRD countries [75]. Coincidentally, these countries receive the least security & privacy research attention, some of which appear in our dataset but have received no prior privacy-focused academic interest [31], making our work instrumental in addressing this gap. We also note that, recent work has not found differences between Android and iOS users’ privacy sensitivities [1], suggesting our results may provide hints beyond Android users. We leave the exploration of this hypothesis to future work.
@@ 5.1 Themes Decreasing in Prevalence
We document here, for the first time, the dominant emotions associated with privacy reviews about app permissions. Unsurprisingly, we find that negative emotions (‘Anger’, ‘Annoyance’, ‘Sadness’, ‘Disgust’, and ‘Fear’), account for 45–68% of reviews for each specific theme (see Figure 4). Qualitative analysis shows that ‘Anger’ reviews commonly question the use case for privacy-sensitive permissions and express disapproval, such as: “Horrible! it doesn’t make sense to give you permission for management of my email account, contacts, and images. All bad ideas.” Among our top 20 themes, the ‘Confusion’ emotion appears most often with privacy-sensitive permission themes (7-14.4%). The permission-related reviews associated with ‘Confusion’ almost exclusively question the need for privacy-sensitive permissions but are less confrontational (e.g., (“Why do you need identity and call information permissions?”).
@@
Other themes exhibiting significant decline in PTPR are App Locking, Content Hiding and Password Protection. These themes collectively have dropped from making up roughly 18% of privacy reviews in Feb 2013 to 5% as of Jan 2023.
We looked to see which app types these reviews come from, and found that the app types were predominantly App lockers/hiders, Album management, and Personal diary & journal. Such apps typically allow a user to protect private content via additional password controls that are especially useful when a device is shared. South Asian countries contribute the most reviews on these themes (e.g., India makes up 29.4% of Content Hiding reviews), and device sharing is a known challenge in these countries [79]). We speculate the decline may be tied to increasing smartphone adoption—and thus reduced device sharing—among south Asian countries [9]. However, more research is needed to confirm this. We further explore these themes and app types in § 7.2.
@@ 5.2 Themes Increasing in Prevalence
We next look at the top four themes that have shown statistically significant increases over time: Data Deletion, Data Theft, Data Usage Concerns, and Unwanted Data Collection. These four themes have collectively risen from 10% of privacy reviews in early Feb 2013 to 27.3% by Jan 2023. Generally speaking, the focus across these themes is on PII and user-generated personal data. We note that our clustering and theme-generation methods are able to distinguish between PII and privacy-sensitive permission discussions.
Themes in this group capture user feedback in various nuanced ways. Data Deletion-themed reviews are almost exclusively requests to delete personal data (discussed in detail in § 5.3). In contrast, Data Theft-themed reviews express concern about data being stolen or an account getting hacked. Data Usage Concerns primarily refers to sharing data with third parties, and Unwanted Data Collection reviews express concerns with the app requesting unnecessary data (e.g., for showing ads).
Overall these reviews occur in two styles. Many reference very specific data types: national IDs (e.g., SSN, TCKN, NIK), financial data (e.g., credit card numbers), personally identifiable information (e.g., email, phone number), or user generated/uploaded content (e.g., “Requires EVERYTHING about you including your social security number in order to scan lottery tickets. Why wouldn’t you use your local store instead of exposing yourself to a risk of data breaches?”). Other reviews associated with these themes discuss personal data vaguely, without reference to a specific type of personal data (e.g., “Warning! The app collects a lot of data. Personal data, device and usage-related data…The app then sells it”).
@@
Finding 3:xxOver time, an increasing share of privacy reviews focus on personal information (e.g., PII, user data). Our analysis is able to differentiate privacy discussions about different types of personal data, such as privacy-sensitive permissions vs PII.
@@ 5.3 The rise in Data Deletion and Data Theft
Since Data Deletion and Data Theft are our 2 largest themes, and both show a statistically significant increase in volume over time, we now look more closely at what may be contributing to this. Our analysis enables us to break down our data along multiple dimensions, thus we looked at per-country contributions for these two topics and found that Indonesia contributes disproportionately. It contributes 40% of the reviews in these two categories, whereas all other countries contribute less than 8% each. We also saw that Indonesia has the fastest PPR growth of any country, with 76% relative growth every two years and the third-most privacy reviews of all countries by volume. Considering only the Data Deletion and Data Theft themes for Indonesia, we next considered app type. We uncovered that 44% of these reviews came from financial loan apps, with other app types contributing less than 0.8% each.
Many Data Theft reviews complain that the app collects PII (social media accounts, photos of IDs), and then either loans are denied without reason, or threats are made leveraging their PII. One reviewer writes ”The app stole my data, including my facebook account. It told me that the loan process would be easier if I input all of my data. My loan was still denied even though I have a good payment history. Watch out people.” Another reviewer says ”[COMPANY] pretends you’ll get a loan with no hiccups. They say they just need your photo identity. If you’re even a day late, their debt collector will harass you and threaten to leak your data (photo identity) to social media.
@@ 6.1 Clustering Countries
We now use these topic distributions to cluster countries. Specifically, we represent each country as a vector of size 50, where each element is the PTPR for a (country, theme) tuple, for the top 50 themes across the entire dataset. We then apply hierarchical clustering with complete linkage [58]. The resulting clusters are shown on a world map (Figure 8). A detailed breakdown can be seen in Figure 12 in Appendix C.
We see from this map that the question of whether geographically close countries write about the same privacy issues yields mixed results. On one hand, Russia, Ukraine, Belarus and Kazakhstan all cluster together (teal). We also see that most of the Anglo countries, namely the U.S., Canada, Australia and the U.K., are in the same (light orange) cluster. However, this light orange cluster also contains Mexico, Argentina, Spain, Portugal, and France. Interestingly, European countries divide between two clusters, with eastern European countries, Germany, and Italy in their own cluster (pink). Similarly, the Middle East and North Africa splits across three main clusters (Türkiye is alone, Iran and Iraq cluster together (yellow), and Saudi Arabia, Algeria, and Egypt cluster together (green). A detailed examination of our dendrogram shows that the Iran-Iraq cluster is surprisingly far away from the Saudi Arabia-Algeria-Egypt cluster. Asian countries exhibit wide diversity. India clusters with some African countries (South Africa, Kenya, and Nigeria), and not with Pakistan or Bangladesh. Interestingly, we note that Indonesia, Pakistan, Thailand and Türkiye appear unique (cluster size 1).
@@ 6.3 Türkiye
Türkiye stands out as anomalous based on two metrics. First, as mentioned above, its distribution of privacy topics discussed does not cluster with any other country (cluster size 1). Second, Türkiye also exhibits anomalous PPR growth over ten years. The middle column of Figure 5 shows the average 2-year change for countries, among the top 20, with statistically significant growth or decline in PPR (KPSS $p\leq 0.05$); stationary (non-changing) countries are excluded. Türkiye shows an unusually large increase, with an average 2-year relative growth of 62.1% between Feb. 2015 and Jan. 2023. (Figure 6 illustrates how its PPR has evolved over time.) Türkiye is also interesting because it provides the 6th-most privacy reviews (502K) of all countries.
Figure 7 shows that Türkiye has an unusual distribution of privacy themes; Fingerprint Matters, Personal Information Privacy and Unneeded Camera Access appear more frequently than in other countries. The privacy reviews highlight user concerns around fingerprint collection (e.g., “My fingerprint was scanned on [date], the company has the responsibility of my fingerprint.”) and abuse (e.g., “If my fingerprint is used in something illegal, the app is responsible.”). Although biometrics such as fingerprints are not shared with apps directly [40],
reviewers express concern about the potential for such sharing, e.g., when a banking app verifies authorization via a fingerprint.

### 15 Mar 2024 v3

@@ Abstract
We present an analysis of 12 million instances of privacy-relevant reviews publicly visible on the Google Play Store that span a 10 year period. By leveraging state of the art NLP techniques, we examine what users have been writing about privacy along multiple dimensions: time, countries, app types, diverse privacy topics, and even across a spectrum of emotions. We find consistent growth of privacy-relevant reviews, and explore topics that are trending (such as Data Deletion and Data Theft), as well as those on the decline (such as privacy-relevant reviews on sensitive permissions). We find that although privacy reviews come from more than 200 countries, 33 countries provide 90% of privacy reviews. We conduct a comparison across countries by examining the distribution of privacy topics a country’s users write about, and find that geographic proximity is not a reliable indicator that nearby countries have similar privacy perspectives. We uncover some countries with unique patterns and explore those herein. Surprisingly, we uncover that it is not uncommon for reviews that discuss privacy to be positive (32%); many users express pleasure about privacy features within apps or privacy-focused apps. We also uncover some unexpected behaviors, such as the use of reviews to deliver privacy disclaimers to developers. Finally, we demonstrate the value of analyzing app reviews with our approach as a complement to existing methods for understanding users’ perspectives about privacy.
@@ 1 Introduction
Getting input about users’ perspectives on privacy issues from millions of user reviews from hundreds of countries has not been feasible until recently. Advances in natural language processing (NLP) and large language models (LLMs) permit complex analysis of enormous text corpora. Here, we leverage these advances to investigate users’ privacy opinions and concerns from a novel perspective: we study 12.3M (million) privacy-related reviews, extracted from $\sim$2B (billion) public reviews on Google Play spanning 10 years (Jan. 2013–Feb. 2023). These privacy reviews come from more than 200 countries or regions, 25 languages, and 160K (thousand) apps that span every Play app category.
@@ 3.1 Text Analysis Pipeline
We build our analysis pipeline to leverage and extend the components in Hark, an end-to-end system for retrieval and analysis of privacy-related feedback leveraging state-of-the-art techniques in NLP [32]. An overview of our data-analysis pipeline is in Figure 1. First, the privacy classifier identifies the privacy-related feedback from unstructured text. The issue generation model takes in this privacy feedback and dynamically generates meaningful, fine-grained issues (covering both known and newly emerging issues) describing the privacy aspects discussed within each review text. The theme creation component groups these issues into thematic clusters and assigns a succinct title to each. Our data-analysis pipeline also includes an emotion classifier that dissects each review’s text across 28 emotions (e.g., anger, fear, joy and confusion).
@@
We obtained models and training data from Hark [32] and re-purposed most of the pipeline, with modifications to the privacy classifier to improve the breadth of topics identified. First, we expand the privacy taxonomy from the original paper (25 concepts) to include 89 privacy-relevant concepts(see Appendix D). The prior taxonomy already included concepts from multiple known taxonomies. However, when we asked five privacy experts (with at least 5 years of academic/industry experience in privacy research and development) to manually examine the original taxonomy [32], they observed many missing topics. For example, we added Data Hiding, Opt-out and Location concepts, and added specificity to the Anonymity and Personal Data types.
@@
Our classifier has a 0.95 ROC AUC, 87% precision, and 86% recall on our new diversified test set, [^1] We set a 0.8 prediction score threshold for best accuracy and verified results qualitatively. and a similar ROC AUC of 0.88 (vs 0.92) on the original Hark test set [32]. In comparison, the privacy classifier from [32] has 0.87 ROC AUC, 89% precision, and 51% recall on our new test set, demonstrating that our classifier performs better at capturing the diversity of our taxonomy. (For additional experiments with a range of model architectures see Appendix B.) We also performed qualitative assessment of all 50 false positives/negatives of our classifier, noticing primarily issues with reviews that were short ($\leq$ 10 words) and ambiguous (missing context, causing multiple possible interpretations). For example, a short review “Taking too much data” was labeled as ‘privacy,’ when the privacy expert (based on other reviews seen) interpreted it to be a complaint about the app consuming the limited mobile data bandwidth and annotated it as ‘not-privacy’; whereas an ambiguous review “msgs have been deleted, but for some reason they remain in place” was labeled as ‘not-privacy’ (could be seen as a bug in app functionality), when the privacy expert interpreted it as data deletion control not working as expected (a privacy concern). 13 of the 50 false positives/negatives were short; and the rest (though longer) were still ambiguous. Despite missing these short and ambiguous cases, our classifier had good performance.
@@ 3.2 Data Description
We obtained access to reviews directly from the Google Play team. Our initial dataset consists of all $\sim$2B publicly visible Google Play app reviews for apps with $>$10K installs, spanning 10 years from Jan 2013–Feb 2023. The dataset was already anonymized (no user identifying information) and sanitized (any detected fake/spam reviews were removed). Each review is associated with the review text, its language, country, submission time, star rating, the corresponding app’s package name, the app’s developer-specified category information and (if available) finer-grained app-type information (such as ‘Rideshare & Taxis’ for a rideshare app, which belongs to the ‘Maps & Navigation’ developer-specified category). Note that we do not generally use developer-specified categories as they are too broad [72]. Instead we use the more specific app types that are displayed on Play when viewing apps.
In our dataset, $\sim$65of reviews are non-English. Since our classifiers work on English texts, we leverage Google’s Translation API to translate reviews in 24 non-English languages to English (see Appendix A for the full list) [4]. The final set of English reviews (including translations) constitutes 98% of the initial dataset and contains 1.9B reviews for 160K apps. These reviews come from all developer-specified categories and 445 app types, with representation from more than 200 countries and territories in the world.
@@ 3.5 Ethical considerations
Our work uncovered reviews that contained ads for spying services. We disclosed our findings to Google Play and they removed these harmful reviews before publication.
@@ 3.6 Limitations
Our work focuses on Android users; however, other sizeable platforms exist, raising generalizability concerns depending on the ratio of android users in a country. We note that Android is the most popular mobile operating system in the world and tends to be even more popular in non-WEIRD countries.[82]. Coincidentally, these countries receive the least security & privacy research attention, some of which appear in our dataset but have received no prior privacy-focused academic interest [33], making our work instrumental in addressing this gap. We also note that, recent work has not found differences between Android and iOS users’ privacy sensitivities [1], suggesting our results may provide hints beyond Android users. We leave the exploration of this hypothesis to future work.
@@ 5.1 Themes Decreasing in Prevalence
We document here, for the first time, the dominant emotions associated with privacy reviews about app permissions. Unsurprisingly, we find that negative emotions (‘Anger’, ‘Annoyance’, ‘Sadness’, ‘Disgust’, and ‘Fear’), account for 45–68% of reviews most themes. (see Figure 5). Qualitative analysis shows that ‘Anger’ reviews commonly question the use case for privacy-sensitive permissions and express disapproval, such as: “Horrible! it doesn’t make sense to give permission for management of my email account, contacts, and images. All bad ideas.” Among our top 20 themes, the ‘Confusion’ emotion appears most often with privacy-sensitive permission themes (7-14.4%). The permission-related reviews associated with ‘Confusion’ almost exclusively question privacy-sensitive permissions but are less confrontational (e.g., “Why do you need identity and call information permissions?”).
@@
Other themes exhibiting significant decline in PTPR are App Locking, Content Hiding and Password Protection. These themes collectively have dropped from roughly 18% of privacy reviews in Feb 2013 to 5% as of Jan 2023.
We investigated the app types of these reviews, and found the reviews were predominantly for App lockers/hiders, Album management, and Personal diary & journal. Such apps typically allow protection of private content via password controls that are especially useful when a device is shared. South Asian countries contribute the most to these themes (e.g., India makes up 29.4% of Content Hiding reviews), and device sharing is a known challenge in these countries [86]). We speculate increasing smartphone adoption—and thus reduced device sharing—might explain the decline among south Asian countries [9]. However, more research is needed to confirm. We further explore these themes and app types in § 7.2.
@@ 5.2 Themes Increasing in Prevalence
The top four themes that have shown statistically significant increases over time are: Data Deletion, Data Theft, Data Usage Concerns, and Unwanted Data Collection. These themes have risen in aggregate from 10% of privacy reviews in early Feb 2013 to 27.3% by Jan 2023. Collectively, these themes focus on PII and user-generated personal data. Our analysis methods discern the nuanced difference between PII and privacy-sensitive permission discussions.
Themes in this group capture feedback in nuanced ways. Data Deletion-themed reviews are almost exclusively requests to delete personal data (discussed in § 5.3). In contrast, Data Theft-themed reviews express concern about data being stolen or an account hacked. Data Usage Concerns primarily refers to sharing data with third parties, and Unwanted Data Collection reviews express concerns with the app requesting unnecessary data (e.g., for showing ads).
These reviews occur in two styles. Many reference precise data types: national IDs (e.g., SSN, TCKN), financial data (e.g., credit cards), personally identifiable information (e.g., emails), or user-generated content (e.g., “Requires EVERYTHING about you including your social security number to scan lottery tickets. Why wouldn’t you use your local store instead of exposing yourself to a risk of data breaches?”). Other reviews in these themes discuss personal data vaguely, without reference to specific personal data types (e.g., “Warning! The app collects a lot of data. Personal data, device and usage-related data…The app then sells it”).
@@
Finding 3:xxOver time, an increasing share of privacy reviews focus on personal information (e.g., PII). Our analysis uncovers privacy discussions about different types of personal data, such as privacy-sensitive permissions vs PII.
@@ 5.3 The rise in Data Deletion and Data Theft
Since Data Deletion and Data Theft are our two largest themes, and both show a statistically significant increase in volume over time, we now look at what the contributing factors may be. Our analysis enables us to break down our data along multiple dimensions, thus we looked at per-country contributions for these two topics and found that Indonesia (disproportionately) contributes 40% of the reviews in these two categories, whereas all other countries contribute less than 8% each. We also saw that Indonesia has the fastest PPR growth of any country, with 76% relative biannual growth and the third-most privacy reviews of all countries by volume. Considering only the Data Deletion and Data Theft themes for Indonesia, we next considered app types. We uncovered that 44% of these reviews came from financial loan apps, with other app types contributing less than 0.8% each.
Many Data Theft reviews complain that the app collects PII (social media accounts, photos of IDs), and then either loans are denied without reason, or threats are made leveraging their PII. One reviewer writes “The app stole my data, including my facebook account. It told me that the loan process would be easier if I input all of my data. My loan was still denied even though I have a good payment history. Watch out people.” Another reviewer says “[COMPANY] pretends you’ll get a loan with no hiccups. They say they just need your photo identity. If you’re even a day late, their debt collector will harass you and threaten to leak your data (photo identity) to social media.
@@ 6.1 Clustering Countries
We now use these topic distributions to cluster countries. Specifically, we represent each country as a vector of size 50, where each element is the PTPR for a (country, theme) tuple, for the top 50 themes across the entire dataset. We then apply hierarchical clustering with complete linkage [61]. The resulting clusters are shown on a world map in Figure 8. A detailed breakdown can be seen in Figure 12 in Appendix C.
We see from this map that the question of whether geographically close countries write about the same privacy issues yields mixed results. On one hand, Russia, Ukraine, Belarus and Kazakhstan all cluster together (teal). We also see that most of the Anglo countries, namely the U.S., Canada, Australia and the U.K., are in the same (light orange) cluster. However, this light orange cluster also contains Mexico, Argentina, Spain, Portugal, and France. Interestingly, European countries divide between two clusters, with eastern European countries, Germany, and Italy in their own cluster (pink). Similarly, the Middle East and North Africa splits across three main clusters: Türkiye is alone, Iran and Iraq cluster together (yellow), and Saudi Arabia, Algeria, and Egypt cluster together (green). A detailed examination of our dendrogram shows that the Iran-Iraq cluster is surprisingly far away from the Saudi Arabia-Algeria-Egypt cluster. Asian countries exhibit wide diversity. India clusters with some African countries (South Africa, Kenya, and Nigeria), and not with Pakistan or Bangladesh. Interestingly, we note that Indonesia, Pakistan, Thailand and Türkiye appear unique (cluster size 1).
@@ 6.3 Türkiye
Türkiye stands out as anomalous based on two metrics. First, its distribution of privacy topics does not cluster with any other country (cluster size 1). Second, Türkiye exhibits anomalous PPR growth over ten years. The middle column of Figure 4 shows the average 2-year change for countries, among the top 20, with statistically significant growth or decline in PPR (KPSS $p\leq 0.05$); stationary (non-changing) countries are excluded. Türkiye shows an unusually large increase, with an average 2-year relative growth of 62.1% between Feb. 2015 and Jan. 2023 (Figure 6 illustrates this). Türkiye also provides the 6th-most privacy reviews (502K) of all countries.
Figure 7 shows that Türkiye has an unusual distribution of privacy themes; Fingerprint Matters, Personal Information Privacy and Unneeded Camera Access appear more frequently than in other countries. The privacy reviews highlight user concerns around fingerprint collection (e.g., “My fingerprint was scanned on [date], the company has the responsibility of my fingerprint.”) and abuse (e.g., “If my fingerprint is used in something illegal, the app is responsible.”). Although biometrics such as fingerprints are not shared with apps directly[42], reviewers express concern about the potential for such sharing, e.g., when a banking app verifies authorization via a fingerprint.